An element with mass 730 grams decays by 28.8% per minute. How much of the element is remaining after 10 minutes, to the nearest 10th of a gram?





1
Expert's answer
2022-02-27T16:56:58-0500

Amount remaining=mo(1-28.8÷100)10

=730(1-0.288)10

=730×0.71210

=24.44

=24.4g